OpenStudy (anonymous):

MEDAL and FAN Identify the vertex for the graph of y = -2x2 - 12x + 1.

OpenStudy (anonymous):

(-3,19) (-3, 55) (3, -57) (3, -47)

OpenStudy (joannablackwelder):

When a function is in this form, the easiest way to find the vertex is to use the formula x-coordinate of vertex is -b/2a

OpenStudy (joannablackwelder):

Find the x-coordinate, then substitute to find the y-coordinate.

OpenStudy (joannablackwelder):

Can you identify a, b, and c?

OpenStudy (anonymous):

A= -2x^2 B= -12x C= 1?

OpenStudy (joannablackwelder):

Close. Just the coefficients.

OpenStudy (joannablackwelder):

So, A=-2

OpenStudy (joannablackwelder):

B=-12

OpenStudy (joannablackwelder):

and C=1

OpenStudy (joannablackwelder):

Make sense?

OpenStudy (anonymous):

Yes! Sorry I lost connection @JoannaBlackwelder

OpenStudy (anonymous):

so -b/2a = 12/2*-2= 12/-4= -3

OpenStudy (joannablackwelder):

Yep, awesome! So, that is the x-coordinate. Can you sub in -3 for x and solve for y?